OpenStudy (anonymous):

A couple needed to purchase a present for 14 members of their families. If the husband bought x presents, how many presents did the wife need to buy?

OpenStudy (anonymous):

Can someone help!

OpenStudy (the_fizicx99):

Wouldn't this be y = 14 - x or x + y = 14

OpenStudy (anonymous):

I am not sure

OpenStudy (the_fizicx99):

do you have options?

OpenStudy (anonymous):

No its just this question

OpenStudy (anonymous):

I first thought the answer was 14x

OpenStudy (the_fizicx99):

@mathmale I think this would be y = 14 - x By subtracting how many the father bought then you can determine how many she needs to buy.

OpenStudy (anonymous):

ok

OpenStudy (the_fizicx99):

@bibby

OpenStudy (anonymous):

Thanks for your help